5. Complete the table below that calculates the Future Value of $10,000 invested at 7% per year using periods from 1 to 10 years. The table also illustrates the value today of $10,000 to be received in the future from 1 to 10 years later using a discount rate of 5%. (Note: table is to be completed by appropriately copying the Excel formulas. Using Excel copy/paste functionality it should take just a few seconds to do it).

Number of

Number of

Number of

Number of

Number of

Number of

Periods

(Excel Formula)

(Algebra Formula)

(Excel Formula)

(Algebra Formula)

1

$10,700.00

$10,700.00

$9,523.81

$9,523.81

2

$11,449.00

$11,449.00

$9,070.29

3

4

5

6

7

8

9

10

$6,139.13

$6,139.13
Answer #1
No.of Pds. (Excel Formula) (Algebra Formula) No.of Pds. (Excel Formula) (Algebra Formula)
1 FV(7%,1,,10000)= 10700 10000*(1+0.07)^1= 10700 1 PV(5%,10,,10000)= $6,139.13 10000/1.05^10= 6139.13
2 FV(7%,2,,10000)= 11449 10000*(1+0.07)^2= 11449 2 PV(5%,9,,10000)= $6,446.09 10000/1.05^9= 6446.09
3 FV(7%,3,,10000)= 12250.43 10000*(1+0.07)^3= 12250.43 3 PV(5%,8,,10000)= $6,768.39 10000/1.05^8= 6768.39
4 FV(7%,4,,10000)= 13107.96 10000*(1+0.07)^4= 13107.96 4 PV(5%,7,,10000)= $7,106.81 10000/1.05^7= 7106.81
5 FV(7%,5,,10000)= 14025.52 10000*(1+0.07)^5= 14025.52 5 PV(6%,6,,10000)= $7,049.61 10000/1.06^6= 7049.61
6 FV(7%,6,,10000)= 15007.30 10000*(1+0.07)^6= 15007.30 6 PV(5%,5,,10000)= $7,835.26 10000/1.05^5= 7835.26
7 FV(7%,7,,10000)= 16057.81 10000*(1+0.07)^7= 16057.81 7 PV(5%,4,,10000)= $8,227.02 10000/1.05^5= 8227.02
8 FV(7%,8,,10000)= 17181.86 10000*(1+0.07)^8= 17181.86 8 PV(5%,3,,10000)= $8,638.38 10000/1.05^3= 8638.38
9 FV(7%,9,,10000)= 18384.59 10000*(1+0.07)^9= 18384.59 9 PV(5%,2,,10000)= $9,070.29 10000/1.05^2= 9070.29
10 FV(7%,10,,10000)= 19671.51 10000*(1+0.07)^10= 19671.51 10 PV(5%,1,,10000)= $9,523.81 10000/1.05^1= 9523.81
Note: While using EXCEL function, you add NEGATIVE sign to the current investment ,( ie. -10000 ), so that the final answer comes out POSITIVE-- as the intial investment is a cash outflow.
EXCEL Nos. C D E F G H I J K
No.of Pds. (Excel Formula) (Algebra Formula) (Excel Formula) (Algebra Formula)
35 1 FV($D$49,C35,0,-$E$49,0)= 10700 $E$49*(1 + $D$49)^C35= 10700 PV($G$49,C35,0,-$H$49,0)= 9523.81 PV($G$49,C35,0,-$H$49,0)= 9523.81
36 2 FV($D$49,C36,0,-$E$49,0)= 11449 $E$49*(1 + $D$49)^C36= 11449 PV($G$49,C36,0,-$H$49,0)= 9070.29 PV($G$49,C36,0,-$H$49,0)= 9070.29
37 3 FV($D$49,C37,0,-$E$49,0)= 12250.43 $E$49*(1 + $D$49)^C37= 12250.43 PV($G$49,C37,0,-$H$49,0)= 8638.38 PV($G$49,C37,0,-$H$49,0)= 8638.38
38 4 FV($D$49,C38,0,-$E$49,0)= 13107.96 $E$49*(1 + $D$49)^C38= 13107.96 PV($G$49,C38,0,-$H$49,0)= 8227.02 PV($G$49,C38,0,-$H$49,0)= 8227.02
39 5 FV($D$49,C39,0,-$E$49,0)= 14025.52 $E$49*(1 + $D$49)^C39= 14025.52 PV($G$49,C39,0,-$H$49,0)= 7835.26 PV($G$49,C39,0,-$H$49,0)= 7835.26
40 6 FV($D$49,C40,0,-$E$49,0)= 15007.30 $E$49*(1 + $D$49)^C40= 15007.30 PV($G$49,C40,0,-$H$49,0)= 7462.15 PV($G$49,C40,0,-$H$49,0)= 7462.15
41 7 FV($D$49,C41,0,-$E$49,0)= 16057.81 $E$49*(1 + $D$49)^C41= 16057.81 PV($G$49,C41,0,-$H$49,0)= 7106.81 PV($G$49,C41,0,-$H$49,0)= 7106.81
42 8 FV($D$49,C42,0,-$E$49,0)= 17181.86 $E$49*(1 + $D$49)^C42= 17181.86 PV($G$49,C42,0,-$H$49,0)= 6768.39 PV($G$49,C42,0,-$H$49,0)= 6768.39
43 9 FV($D$49,C43,0,-$E$49,0)= 18384.59 $E$49*(1 + $D$49)^C43= 18384.59 PV($G$49,C43,0,-$H$49,0)= 6446.09 PV($G$49,C43,0,-$H$49,0)= 6446.09
44 10 FV($D$49,C44,0,-$E$49,0)= 19671.51 $E$49*(1 + $D$49)^C44= 19671.51 PV($G$49,C44,0,-$H$49,0)= 6139.13 PV($G$49,C44,0,-$H$49,0)= 6139.13
49 7% 10000 5% 10000